OpenStudy (anonymous):

This is probability. Can someone please help? In a manufacturing process, 0.2% of the brake pads that come off the production line are considered defective. Determine the probability that at least one of the brake pad is defective in an order of 200 pads.

OpenStudy (kropot72):

\[P(0\ defective)=\left(\begin{matrix}200 \\ 0\end{matrix}\right)0.002^{0}\times0.998^{200}=0.998^{200}\] \[P(at\ least\ 1\ is\ defective)=1.000-0.998^{200}=you\ can\ calculate\]